OpenStudy (anonymous):

Given the function of f(x)=4(x+1)^2-3, indicate the shifts that will effect the locations of the vertex and explain what effect they will have. Use complete sentences f(x-2) f(x)-2 f(2x) 2f(x) @hartnn @SolomonZelman @dan815 @AravindG @Compassionate @iGreen

OpenStudy (dan815):

the x+1 makes it go 1 to the left and the -3 makes it go down 3

OpenStudy (dan815):

for f(x-2) it will shift your function 2 to the right f(X)-2 all of the function is shifted down by 2 f(2x) makes it happen twice as fast so its compressed horizontally by 2 2(f(x)) makes all your y calues 2 times bigger so its stretched vertically by 2

OpenStudy (dan815):

for example

OpenStudy (dan815):

here is a way to picture it

OpenStudy (dan815):

suppose f(x)=x^2 okay now f(x-2)=(x-2)^2 we know x^2 is a parabola and -2 will maek the parabola shift to the right by 2 but another way to see why it does shift by 2 lets change our axis f(x-2)=f(a)=a^2 if a=x-2 then a+2=x if u graph a,f(a) its just a parabola originating from (0,0)as its f(a)=a^2 but we know this means in terms of x its a+2 so the x version is shifted 2 to the right